# Multi-Agent Causal Discovery Using Large Language Models

### v1

@@ 1 Introduction
The advent of Large Language Models (LLMs), trained on vast amounts of data, has enabled them to acquire extensive knowledge, from common sense to specific domains such as math and science. Recent studies suggest that complex behaviors, such as writing code, generating long stories, and even reasoning capabilities, can emerge from large-scale training (Wei et al. (2023); Rozière et al. (2024); Zhao et al. (2023b); Yao et al. (2023a)). LLMs present a promising alternative for obtaining expert knowledge more accessible and affordable. Recent research (Kıcıman et al. (2023);Choi et al. (2022); Long et al. (2024)) has attempted to leverage these capabilities for causal discovery based on metadata and knowledge-based reasoning, akin to human domain experts.
Figure 1: Meta-Debate Module
@@
The first is called Meta Agents Model, which includes a Meta-Debate Module with two debater agents and one judge agent. These agents are designed for causal discovery problems and utilize the nature of debating to find causal relationships among variables through multiple rounds of discussion. Given a list of variables in tabular form, this group determines the causal relationships and outputs the causal graph in matrix form.
The second is called Coding Agents Model, which includes a Debate-Coding Module with four agents operating in two phases. This group leverages both the debating abilities of the agents and statistical causal algorithms. In the first phase, a meta-question for determining a suitable statistical causal discovery algorithm for a particular data setting is presented to the Meta-Debate Module. After the debate, an algorithm is chosen, and the winning agent provides a step-by-step plan. In the second phase, this plan is then given to a causal coding executor agent, who writes, executes, and debugs the code to achieve a matrix representation of the causal graph.
The third is called Hybrid Model, which hybridizes the statistical causal discovery algorithms with the reasoning skills of multiple agents to construct the causal graph. There are two hybrid mechanisms. The first is the Coding-Meta Hybrid. It includes two phases, for the first phase uses a similar approach of the Coding Agents Model and then leverages this output as an additional parameter for the second phase which uses the Meta Agents Model. Conversely, the second is the Debating-Coding Hybrid. It implements the Meta Agent Model first, then uses the result as an additional parameter for the Coding Agents Model. The final output of either of the two Hybrid Models also yields a matrix representation of the causal graph.
@@ 2 Related Works
2.1 Prompt Engineering
Several advanced techniques for leveraging large language models (LLMs) have been identified. Zero-shot prompting, introduced by Radford et al. (2019), guides LLMs to perform novel tasks using carefully crafted prompts without the need for training data, allowing the model to leverage its existing knowledge to generate predictions. Few-shot prompting, as described by Brown et al. (2020), enhances model performance by providing a few input-output examples, though it requires more tokens and careful example selection to mitigate biases. For reasoning and logic, Wei et al. (2023) introduced Chain-of-Thought (CoT) prompting, which guides LLMs through step-by-step reasoning processes, significantly improving accuracy in complex tasks such as math and commonsense reasoning. Building on this, Wang et al. (2023) proposed self-consistency, a strategy that generates diverse reasoning chains and identifies the most consistent final answer, further enhancing accuracy. Additionally, Yao et al. (2023b) developed ReAct, enabling LLMs to generate reasoning traces and task-specific actions concurrently, thereby improving performance in question answering, fact verification, and interactive decision-making by enhancing the synergy between reasoning and action. In this work, due to deliberate thinking such as analyzing the problem or code writing, we integrate ReAct as one of our main prompting techniques and also one-shot prompting when the agents are required to output in a certain way.
@@ 2.2 LLMs’ Agentic Workflow
However, for more complex problems where a single LLM agent may struggle, LLM-MA (multi-agent) systems excel. Current LLM-MA systems primarily employ three communication paradigms: Cooperative, Competitive, and Debating. In the Cooperative paradigm, agents collaborate towards a shared goal, typically exchanging information to enhance a collective solution Qian et al. (2023); Chen et al. (2024b). In the Competitive paradigm, agents work towards their own goals, which might conflict with those of other agents Zhao et al. (2023a). The Debating paradigm involves agents engaging in argumentative interactions, where they present and defend their viewpoints or solutions while critiquing those of others. This approach is ideal for reaching a consensus or a more refined solution (Li et al. (2023); Liang et al. (2023); Xiong et al. (2023)). In this work, the debating paradigm will be implemented, as the nature of causal discovery problems requires diverse and potentially conflicting opinions to approach the truth.
@@ 2.3 Statistical and LLM-based Causal Methods
Traditional methods of statistical causal inference often depend heavily on assumed causal graphs to identify and measure causal impacts. To overcome this limitation, data-driven algorithmic approaches have been developed into statistical causal discovery (SCD) methods, encompassing both non-parametric (e.g., Spirtes et al. (2000); Chickering (2002); Silander and Myllymäki (2006); Yuan and Malone (2013); Huang et al. (2018); Xie et al. (2020)) and semi-parametric (e.g., Shimizu et al. (2006); Hoyer et al. (2009); Shimizu et al. (2011); Rolland et al. (2022); Tu et al. (2022)) techniques. Many SCD algorithms can be systematically augmented with background knowledge and have accessible software packages. For example, the non-parametric and constraint-based Peter-Clerk (PC) algorithm (Spirtes et al. (2000)) in “causal-learn” integrates background knowledge of mandatory or forbidden directed edges. “Causal-learn” also includes the Exact Search algorithm (Shimizu et al. (2011); Yuan and Malone (2013) ), a non-parametric and score-based SCD method that can incorporate background knowledge in the form of a super structure matrix of forbidden directed edges. Furthermore, the semi-parametric DirectLiNGAM (Shimizu et al. (2011)) algorithm can use prior knowledge of causal order (Inazumi et al. (2010)) in the “LiNGAM” project (Ikeuchi et al. (2023)).
Figure 2: Debate-Coding Module
In the context of knowledge-driven approaches using large language models (LLMs), applying LLMs for causal inference is relatively new. There have been a few significant efforts to use LLMs for causal inference among variables by merely prompting with the variable names, without going through the traditional SCD process with benchmark datasets (Kıcıman et al. (2023); Zečević et al. (2023)). Jiralerspong et al. (2024) even uses a breadth-first search (BFS) approach which allows it to use only a linear number of queries to have a higher efficiency. Researchers have focused on using LLMs to minimize error metrics in score-based SCD results (Ban et al. (2023)) and to improve causal effect estimation by determining the causal order from triplet subgraphs derived from non-parametric SCD results (Vashishtha et al. (2023)). Another approach focuses on constructing background knowledge based on the response probability of the LLM to reflect the credibility of its decisions with SCP and details achieving statistical validity and natural interpretation in semi-parametric SCD methods like LiNGAM using causal coefficients and bootstrap probabilities (Takayama et al. (2024)).
@@ 3.1.1 Meta-Debate Module
The Meta-Debate Module is an advanced system comprising three intelligent agents: two causal debaters (an affirmative one and a negative one) and one causal judge. This structure emulates the dynamic and rigorous nature of human debate, specifically within the realm of causal discovery. The design of this debating module is meticulously crafted to ensure a thorough examination of critical elements in causal discovery, such as understanding the temporal order necessary to establish cause-and-effect relationships and identifying potential confounding variables that could distort perceived relationships between primary variables. Each side engages in active disagreement by presenting different opinions and viewpoints, fostering a comprehensive and robust debate. Additionally, each agent within Meta-Debate Module utilizes the ReAct prompting technique. This technique integrates the ability to dynamically formulate, modify, and refine action plans based on new information or insights gained during the debate. This integration allows the agents to engage in more sophisticated and adaptive reasoning processes, closely mimicking human-like debate and decision-making 3.
@@
For meta-questions, questions might address aspects such as the appropriate algorithm for a particular problem, the causal relationship between two variables, or a step-by-step approach for solving a causal problem, as illustrated in Figure 1. The affirmative side initiates the debating process based on the constructed input. The final output is derived from the causal judge’s decision, reflecting a comprehensive evaluation of the arguments presented by both sides. This rigorous process ensures that the outcome is well-informed and considers multiple perspectives, thereby enhancing the robustness of causal discovery. A detailed visualization of this module can be found in Appendix A.1.
Figure 3: Prompting design of each agent in MAC, details in the Appendix B
@@ 3.1.2 Debate-Coding Module
In the second phase, the causal coding executor receives the plan from the previous phase along with the observational data. The causal coding executor is responsible for writing, executing, and debugging the code. It pre-prompts the functions and provides parameters within a specific Python library [^1] More functions can be found at https://causal-learn.readthedocs.io/en/latest/index.html based on the algorithm selected by the debaters in the initial stage. This pre-prompting is crucial because LLMs can call functions from their training dataset, which may be outdated or incorrect, leading to errors and excessive debugging (details of prompting design can be found in the Appendix B). After executing the code, the causal coding executor outputs a matrix representation of the causal graph. A detailed visualization of this module can be found in the Appendix A.2.
@@ 3.2.1 Meta Agents Model
8:         $G_{ij}$leftarrow Meta_Debate_Module("Are there any direct causal from $X[i]$ to $X[j]$")
@@
11:         $G_{ij}$leftarrow Meta_Debate_Module("Are there any direct causal relationship from $X[i]$ to $X[j]$")
@@ 4 Experiment and Results
4.1 Experimental Setup
We use GPT-3.5-turbo API for most of our experiments with the temperature set at 0. We experiment on three different datasets that adopted from the work of Takayama et al. (2024)
  1. 1.
    Auto MPG dataQuinlan (1993) This dataset consists of the variables around the fuel consumption of cars. With five variables: “Weight”, “Displacement”, “Horsepower”, “Acceleration” and “Mpg”(miles per gallon)
  2. 2.
    DWD climate dataMooij et al. (2016). This dataset encompasses six continuous variables capturing climate observations such as altitude, temperature, precipitation levels, longitude, sunshine duration, and latitude. It is aimed at studying weather patterns, climate change impacts, and geographical correlations in climate variables.
  3. 3.
    Sachs protein dataSachs et al. (2005). The dataset comprises protein signaling measurements from multiparameter single-cell data, capturing the interactions among various proteins (raf, mek, plc, pip2, pip3, erk, akt, pka, pkc, p38, jnk). It’s aimed at understanding signal transduction pathways within cells, derived from an influential study published in Science.
4.2 Metrics
For the evaluation metrics, we assess the adjacency matrix obtained from LLMs or a code executor using structural hamming distance (SHD), false positive rate (FPR), false negative rate (FNR), precision, and F1 score, following the methodology of Takayama et al. (2024). Additionally, we use the Normalized Hamming Distance (NHD) as described by Kıcıman et al. (2023).
4.3 Results
This experiment uses several metrics to evaluate different methods for constructing graph adjacency matrices from datasets with continuous variables. We compare with some classical statistical causal discovery algorithms such as PC, Exact Search, and DirectLiNGAM. We also compare with LLMs-based statistical methods from the work of Takayama et al. (2024) (run on the GPT-3.5-turbo first pattern): Peter-Clark (PC) LLM-KBCI, Exact Search (ES) LLM-KBCI, and DirectLiNGAM LLM-KBCI. Lastly, we also compare with single-agent zero-shot Chain-of-Thought with GPT-3.5-turbo and GPT-4o (checkpoint on 13 May 2024).
@@ 4.3 Results
For the Auto MPG data, the Coding Agents Model achieves the best performance in terms of SHD (4), NHD (0.48), FNR (0), Precision (1), Recall (0.357), and F1 score (0.526). The Meta Agents Model also performs well with an SHD of 5, NHD of 0.2, FNR of 0.4, FPR of 0.15, Precision of 0.5, Recall of 0.6, and F1 score of 0.545. The traditional methods and other advanced methods show higher SHD values, indicating lower structural accuracy. For example, PC, DirectLiNGAM, and Single-agent zero-shot prompting (GPT-4o) all have SHD values of 8, demonstrating less accurate structural learning compared to the Coding agents.

### v2

@@ 1 Introduction
The advent of Large Language Models (LLMs), trained on vast amounts of data, has enabled them to acquire extensive knowledge, from common sense to specific domains such as math and science. Recent studies suggest that complex behaviors, such as writing code, generating long stories, and even reasoning capabilities, can emerge from large-scale training (Wei et al. (2023); Rozière et al. (2024); Zhao et al. (2023b); Yao et al. (2023a)). LLMs present a promising alternative for obtaining expert knowledge more accessible and affordable. Recent research (Kıcıman et al. (2023);Choi et al. (2022); Long et al. (2024); Chen et al. (2024b)) has attempted to leverage these capabilities for causal discovery based on metadata and knowledge-based reasoning, akin to human domain experts.
@@
Method / Approach LLMs for metadata Statistical Approach for structured data Agentic Ability Multi-Agent Workflow Introduced By
Pairwise Causal Discovery × × × Kıcıman et al. (2023); Zečević et al. (2023)
Various of prompt-engineering strategies × × × Chen et al. (2024b)
Efficiently asking question for Causal Graph Discovery Using LLMs × × Jiralerspong et al. (2024)
Combining LLMs with traditional causal methods. × × Vashishtha et al. (2023); Takayama et al. (2024)
Multi-Agent Causal Discover (MAC) Our approach
Table 1: Comparison of Approaches for Using LLMs in Causal Discovery
The first is called Meta Agents Model, which includes a Meta-Debate Module with two debater agents and one judge agent. These agents are designed for causal discovery problems and utilize the nature of debating to find causal relationships among variables through multiple rounds of discussion. The second is called Coding Agents Model, which includes a Debate-Coding Module with four agents operating in two phases. This group leverages both the debating abilities of the agents and statistical causal algorithms. The third one is Hybrid Model, which hybridizes the statistical causal discovery algorithms with the reasoning skills of multiple agents to construct the causal graph.
In this research, we experiment with various LLM models across small, medium, and large scales, paired with different MAC models. We conduct an in-depth analysis of performance for each MAC model and LLM, token consumption, and identify their usage patterns and limitations. Additionally, we propose alternative solutions to address the computational costs associated with LLMs.
@@ 2.1 LLMs’ Agentic Workflow
However, for more complex problems where a single LLM agent may struggle, LLM-MA (multi-agent) systems excel. Current LLM-MA systems primarily employ three communication paradigms: Cooperative, Competitive, and Debating. In the Cooperative paradigm, agents collaborate towards a shared goal, typically exchanging information to enhance a collective solution Qian et al. (2023); Chen et al. (2024c). In the Competitive paradigm, agents work towards their own goals, which might conflict with those of other agents Zhao et al. (2023a). The Debating paradigm involves agents engaging in argumentative interactions, where they present and defend their viewpoints or solutions while critiquing those of others. This approach is ideal for reaching a consensus or a more refined solution (Li et al. (2023); Liang et al. (2023); Xiong et al. (2023)). In this work, the debating paradigm will be implemented, as the nature of causal discovery problems requires diverse and potentially conflicting opinions to approach the truth.
@@ 2.2 Statistical and LLM-based Causal Methods
Traditional methods of statistical causal inference often depend heavily on assumed causal graphs to identify and measure causal impacts. To overcome this limitation, data-driven algorithmic approaches have been developed into statistical causal discovery (SCD) methods, encompassing both non-parametric (e.g., Spirtes et al. (2000); ; ; Yuan & Malone (2013); Huang et al. (2018); Xie et al. (2020)) and semi-parametric (e.g., Shimizu et al. (2006); Hoyer et al. (2009); Shimizu et al. (2011)) techniques. Many SCD algorithms can be systematically augmented with background knowledge and have accessible software packages. For example, the non-parametric and constraint-based Peter-Clerk (PC) algorithm (Spirtes et al. (2000)) in “causal-learn” integrates background knowledge of mandatory or forbidden directed edges. “Causal-learn” also includes the Exact Search algorithm (Shimizu et al. (2011); Yuan & Malone (2013) ), a non-parametric and score-based SCD method that can incorporate background knowledge in the form of a super structure matrix of forbidden directed edges. Furthermore, the semi-parametric DirectLiNGAM (Shimizu et al. (2011)) algorithm can use prior knowledge of causal order (Inazumi et al. (2010)) in the “LiNGAM” project (Ikeuchi et al. (2023)).
In the context of knowledge-driven approaches using large language models (LLMs), applying LLMs for causal inference is relatively new. There have been a few significant efforts to use LLMs for causal inference among variables by merely prompting with the variable names, without going through the traditional SCD process with benchmark datasets (Kıcıman et al. (2023); Zečević et al. (2023)). Jiralerspong et al. (2024) even uses a breadth-first search (BFS) approach which allows it to use only a linear number of queries to have a higher efficiency. Additionally, Chen et al. (2024b) proposes 9 prompting techniques comprise ICL, 0-shot CoT (e.g. ”let’s think step by step”), adversarial prompt, manual CoT, and explicit function (e.g using encouraging language in prompts). Moreover, the works of Ban et al. (2023) and Vashishtha et al. (2023) introduce interesting approach by integrate LLMs into traditional data-driven approaches. However, most of above-mentioned works have not investigate the LLM-agentic work-flows for causal graph discovery 1, which indeed requires heavy investigation on various models and graphs’ scales.
@@ 3.1.1 Meta-Debate Module
Figure 1: Meta-Debate Module
The Meta-Debate Module is an advanced system comprising three intelligent agents: two causal debaters (an affirmative one and a negative one) and one causal judge. This structure emulates the dynamic and rigorous nature of human debate, specifically within the realm of causal discovery. The design of this debating module is meticulously crafted to ensure a thorough examination of critical elements in causal discovery, such as understanding the temporal order necessary to establish cause-and-effect relationships and identifying potential confounding variables that could distort perceived relationships between primary variables. Each side engages in active disagreement by presenting different opinions and viewpoints, fostering a comprehensive and robust debate. Additionally, each agent within Meta-Debate Module utilizes the ReAct prompting technique. This technique integrates the ability to dynamically formulate, modify, and refine action plans based on new information or insights gained during the debate. This integration allows the agents to engage in more sophisticated and adaptive reasoning processes, closely mimicking human-like debate and decision-making (Figure 2).
Figure 2: Prompting design of each agent in MAC, details in the Appendix A.3
@@
For meta-questions, questions might address aspects such as the appropriate algorithm for a particular problem, the causal relationship between two variables, or a step-by-step approach for solving a causal problem, as illustrated in Figure 1. The affirmative side initiates the debating process based on the constructed input. The final output is derived from the causal judge’s decision, reflecting a comprehensive evaluation of the arguments presented by both sides. This rigorous process ensures that the outcome is well-informed and considers multiple perspectives, thereby enhancing the robustness of causal discovery. A detailed visualization of this module can be found in Appendix A.2.1.
Figure 3: Debate-Coding Module
@@ 3.1.2 Debate-Coding Module
In the second phase, the causal coding executor receives the plan from the previous phase along with the observational data. The causal coding executor is responsible for writing, executing, and debugging the code. It pre-prompts the functions and provides parameters within a specific Python library [^1] More functions can be found at https://causal-learn.readthedocs.io/en/latest/index.html based on the algorithm selected by the debaters in the initial stage. This pre-prompting is crucial because LLMs can call functions from their training dataset, which may be outdated or incorrect, leading to errors and excessive debugging (details of prompting design can be found in the Appendix A.3). After executing the code, the causal coding executor outputs a matrix representation of the causal graph. A detailed visualization of this module can be found in the Appendix A.2.2.
@@ 3.2.1 Meta Agents Model
8:         $G_{ij}$leftarrow Meta_Debate_Module(”Are there any direct causal from $X[i]$ to $X[j]$”)
@@
11:         $G_{ij}$leftarrow Meta_Debate_Module(”Are there any direct causal relationship from $X[i]$ to $X[j]$”)
@@ 4 Experiment and Results
4.1 Experimental Setup and Metrics
We use OpenAI API including GPT-3.5, GPT-4o, and GPT-4o mini for our experiment, Groq API for Llama-8.1-70b, Llama-8.1-8b, and Gemini-9B for most of our experiments with the temperature set at 0. We experiment on three different datasets that adopted from the work of Takayama et al. (2024) including Auto MPG data (Quinlan (1993)), DWD climate data (Mooij et al. (2016)), and Sachs protein data (Sachs et al. (2005)). For the evaluation metrics, we assess the adjacency matrix obtained from LLMs or a code executor using structural hamming distance (SHD) and Normalized Hamming Distance (NHD) as described by Takayama et al. (2024) and Kıcıman et al. (2023) respectively.
4.2 Performance of MAC
The results of our experiments indicate that the performance is highly dependent on the complexity of the dataset. We rank and summarize the performance of each model based on the empirical results: (1) Coding Agent, this model exhibits strong performance, particularly when larger models. (2) Coding-Debating Hybrid / Debating-Coding Hybrid, these models provide balanced performance but do not excel in handling highly complex datasets. They are particularly effective in moderate complexity settings. (3) Causal Agent Debate, this model performs well for simpler datasets but faces challenges with more complex ones.
@@ 4.2 Performance of MAC
Auto MPG Dataset: In the Auto MPG dataset, which has a 5x5 matrix graph’s structure, the Causal Agent Debate method performed strongly, particularly with GPT-4o mini, achieving the lowest SHD of 4 and a NHD of 0.16. The Coding Agent method also showed good results, with GPT-3.5 achieving a similarly low SHD of 4 but a higher NHD of 0.48. The Coding-Debating Hybrid method demonstrated balanced performance, with both GPT-4o mini and Llama 3.1 8b achieving an SHD of 5, while Llama 3.1 8b had the lowest NHD of 0.2. The traditional methods and other advanced methods show higher SHD values, indicating lower structural accuracy. For example, PC, DirectLiNGAM, and Single-agent zero-shot prompting (GPT-4o) all have SHD values of 8, demonstrating less accurate structural learning compared to the Coding agents